RHIC pp Set-Up: Over the long weekend there were problems with the microswitches on the door at 4 o'clock. The PASS system brought down RHIC on Sunday at 3:11AM, and initial work on the PLC was not successful. Replacing the switch did not help the first time, since the replacement switch was bad. Replacing the switch again solved the problem, but the PASS system brought RHIC off for over 12 hours. Very frustrating.
The timing of the injection kickers was found to be wrong. Fixing this reduced the emittances in 109 bunch running.
Development of the 80cm b* is expected to continue today.
The JET has accumulated ~1 million events in Blue, and expect to move to Yellow. This should give a polarization measurement at the 8% statistical level.
STAR says Blue and Yellow polarizations are about the same. The STAR local Polarimeters indicate polarization that is consistent with ~50%.
Source polarization is ~80%. Polarization out of the AGS at Flattop is ~60%. We have the tools to measure the polarization profile in both horizontal and vertical in the AGS. We should do this now.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |